# Following Is All You Need: Robot Crowd Navigation Using People As Planners

### v1

@@ I Introduction
Crowd navigation is a challenging problem in robotics research as it involves not only static environmental obstacles, but also dynamic agents such as humans. To handle such complex scenarios, modern planners are designed with techniques such as predictive planning, reinforcement learning, and imitation learning [1]. Recent research focuses on further improving the performance of these advanced planners through more accurate future estimation [2] or more diverse and efficient training [3, 4]. All these works aim to develop intelligent and sophisticated robotic systems that can make correct planning decisions, especially in dense crowds and safety-critical scenarios. On the other hand, humans, perhaps the most intelligent agents in the scene, are solving similar navigation problems as the robot. If we think outside the box, can the robot take advantage of human intelligence to simplify its task?
@@
Inspired by how children follow their parents when walking in crowded streets, we find that following trustworthy leaders is an alternative to planning independently, as illustrated in Fig. 1. We can decompose the crowd navigation problem into high-level decisions (e.g. how to cut through a crowd, whether to overtake the person in front) and low-level short-horizon movements. The former can be solved using the aforementioned intelligent systems or humans, while the latter only requires simple local planners. Therefore, we propose to rethink human-robot relationships in crowd navigation scenarios by handing over high-level decisions to humans. Apart from being dangerous obstacles that the robot needs to avoid, previous works have explored using people as sensors [5, 6, 7, 8] to extend and enrich robot perception. In this work, we use people as intelligent planners that can help the robot solve the more difficult part of the crowd navigation problem.
@@
Our proposed method consists of two steps, leader selection and subgoal selection. We start by evaluating nearby humans based on their observed trajectories. Through a set of rule-based criteria, we aim to identify suitable human leaders who are 1) ahead of and reachable from the robot’s current position, 2) heading towards a similar goal at a desired speed. We then find a subgoal such that it allows the robot to follow the selected leader closely with minimal interruption from other humans. The subgoals are designed to be easily achievable using a simple base planner and we choose Social Force (SF) [9] in our experiments. The leader selection process is performed continuously so the robot can switch between different humans for maximized efficiency. The robot will fall back to the base planner if there is no suitable leader to follow. This usually happens when the scene is sparse and navigation is straightforward. We therefore simplify the task from complex crowd navigation to following suitable human leaders using simple and lightweight local planners, which also enables the robot to demonstrate human-like planning behaviours.
Of relevance to our work is the literature on human-following robots. Many of these works study how to maintain a proper distance from humans in different indoor and outdoor scenarios to ensure human comfort [10, 11, 12]. Others study how to identify a specific user in crowded environments, or with interrupted observations [13, 14, 15]. In these works, a specific user is predefined for the robot to follow without additional context or objectives. Our work is fundamentally different as 1) the robot actively chooses a suitable leader to follow, 2) by following humans, we effectively address the challenge of crowd navigation.
Our contributions are threefold. First, we propose a novel crowd navigation framework that utilises human intelligence to decompose the original problem into straightforward subgoal planning. Second, we design a set of evaluation processes to find subgoals near suitable human leaders, which can be effectively reached using a simple SF planner. Third, we demonstrate through experiments that the proposed People-as-Planner scheme contributes to safe and efficient robot behaviours in crowded and safety-critical scenarios.
@@ II-A Social Robot Navigation
Robot navigation in crowds, or social navigation, has been studied for decades [1, 16]. Traditional model-based methods such as Dynamic Window Approach [17], Social Force [9], Optimal Reciprocal Collision Avoidance [18] are widely adopted in simulations. However, they are less suitable for real-world complex scenarios as they are short-sighted and tend to over-simplify human behaviours [19]. Predictive planning enables the robot to look into the future and avoid potential collisions. They rely heavily on the performance of the prediction module, as inaccurate predictions may lead to accidents and uncertain predictions result in over-conservative robot behaviours [20, 21]. End-to-end learning-based methods train robot policies using large-scale demonstrations [22, 23, 24, 25]. The performance of the trained networks relies on the accuracy of the simulation platform, or the quality of the data collected from real-world experiments [26]. These existing methods, regardless of the complexity, require the robot to plan its path independently. We propose a different approach where the more complex navigation decisions are solved by humans, and the robot is given a simpler task to reach nearby subgoals.
@@ II-B Human-following Robot
Examples of social navigation scenarios include nursing and warehouse assistance, where robots are tasked to follow a target user. Such systems generally consist of modules including human detection, human tracking, and path planning [27], which have been extensively studied by the research community. Recent works focus on more specific tasks in human-following. [10] proposes frontal following where the robot needs to predict the human’s future trajectories and stay ahead of the human. [13] and [14] use camera-Lidar fusion to track the same person under illumination changes and occlusions. [28] additionally incorporates an RF transponder to identify the target human. There are also works that explore new scenarios such as underwater [29] and aerial [30]. We direct the reader to [27] for a comprehensive review. Existing human-following methods assign a hypothetical task to the robot to follow a predefined human target. In our work, the robot’s objective is instead to navigate to its destination by actively selecting and following human leaders.
@@ II-C Human-Robot Relationships
As the robot navigates in crowds, there exist complex relationships between humans and the robot. Apart from considering humans as dynamic obstacles that the robot needs to avoid, [5] first proposes the concept of using people as sensors to benefit from human intelligence. This line of work infers the positions of occluded obstacles from the avoidance behaviours of humans or human drivers [6, 7, 8, 31]. We propose to use people as planners to help the robot make informed navigation decisions. There are a few works that also consider human-following behaviours in robot planning but with different focuses. [32] and [33] consider human-following as one of the candidate policies when generating trajectory samples in a traditional sampling-based planner. [13] and [34] utilise human guidance for pathfinding in mapless environments with low human density. [35] imitates pedestrian behaviours on a narrow sidewalk with limited forms of interactions between humans and the robot. In comparison, our proposed People-as-Planner framework can be applied to various crowded navigation scenarios and is more suitable for real-world robot deployment.
@@ III Methodology
Problem Definition. The state of an agent consists of its position and velocity $\boldsymbol{x}=[\boldsymbol{p}^{\top},\boldsymbol{v}^{\top}]^{\top}\in\mathbb{R}^{4}$. We use subscripts $H,R$ to represent human and robot respectively, and superscript $t$ to represent timesteps. We assign a unique ID $H_{i}$ where $i\in\{1,\dots,n\}$ to the $n$ humans that are within the robot’s observable range. The historical trajectories of all humans can be represented by $\boldsymbol{X}_{H}^{0:t}$. The robot’s goal position is defined as $\boldsymbol{p}_{g}\in\mathds{R}^{2}$. The environmental obstacles within the robot’s observable range are represented by $\mathcal{O}$. In a crowd navigation setting, the robot plan is generated by a planner $\mathcal{P}$, denoted as
@@
The planner $\mathcal{P}$ replans at every step until the robot reaches its goal.
Overview. In our proposed scheme, we first identify a human leader $H_{L}^{t}$ and subsequently define a subgoal $\boldsymbol{p}_{g}^{t}$:
@@
where $f_{\text{leader}}(\cdot)$ and $f_{\text{subgoal}}(\cdot)$ are the leader and subgoal selection processes respectively that will be detailed in Sec. III-A and Sec. III-B. The robot plan is then generated by a base planner $\mathcal{P}$:
@@
and the process repeats until the robot reaches its goal. We assume that the robot dynamics is incorporated into the planner and our method is planner-agnostic. Since each subgoal $\boldsymbol{p}_{g}^{t}$ is much easier to reach than the original goal $\boldsymbol{p}_{g}$, we can use a simple and lightweight local planner as the base planner. The overall framework is illustrated in Fig. 2.
@@ III-A Leader Selection
Figure 3: An example of the reachability check from a robot (red cross) to three of the candidate leaders (green dot). The purple curves enclose its visible region. Human 4 has a larger score than Human 0 because it is closer to the robot. Human 3 has a negative score because it is blocked by Human 2.
Next, we evaluate the reachability of each human. As we aim to use simple base planners, the selected leader needs to be directly accessible without the need for complex obstacle avoidance. Spatio-temporal planning [36, 37] is a feasible method, but at a high computational cost with prediction uncertainties. Instead, we adopt the Line-of-Sight (LoS) distance metric proposed by [38] to define the reachability. As illustrated in Fig. 3, LoS measures the distance from each human to the boundary of the robot’s visible region, which is constructed from the robot’s LiDAR scans. We represent the reachability score as $f_{\text{reach}}$ and only consider Human $i$ as a potential leader if
@@
$S_{\text{head}}=\begin{cases}\frac{\tilde{\boldsymbol{v}}_{i}\cdot\overrightarrow{\boldsymbol{p}_{i}\boldsymbol{p}_{g}}}{\|\tilde{\boldsymbol{v}}_{i}\|\|\overrightarrow{\boldsymbol{p}_{i}\boldsymbol{p}_{g}}\|}&\text{if }\frac{\tilde{\boldsymbol{v}}_{i}\cdot\overrightarrow{\boldsymbol{p}_{i}\boldsymbol{p}_{g}}}{\|\tilde{\boldsymbol{v}}_{i}\|\|\overrightarrow{\boldsymbol{p}_{i}\boldsymbol{p}_{g}}\|}\geq\cos\frac{\pi}{4},\\ -1,&\text{otherwise}.\end{cases}$ (5)
where $\tilde{\boldsymbol{v}}_{i}=\frac{1}{T}\sum_{k=t-T+1}^{t}\boldsymbol{v}_{i}^{k}$.
@@
$S_{\text{vel}}=\begin{cases}\frac{\tilde{v}_{i}-v_{\text{pref}}}{v_{\text{pref}}},&\text{if }\tilde{v}_{i}<v_{\text{pref}},\\ \max\left(0,1-\frac{\tilde{v}_{i}-v_{\text{pref}}}{v_{\text{pref}}}\right),&\text{otherwise}.\end{cases}$ (6)
@@
Third, we compare relative positions to identify humans between the robot and its goal, favouring those closer to the robot:
$S_{\text{pos}}=\begin{cases}\max\left(0,1-\frac{\|\overrightarrow{\boldsymbol{p}_{R}\boldsymbol{p}_{i}}\|}{r}\right),&\text{if }\frac{\overrightarrow{\boldsymbol{p}_{R}\boldsymbol{p}_{i}}\cdot\overrightarrow{\boldsymbol{p}_{R}\boldsymbol{p}_{g}}}{\|\overrightarrow{\boldsymbol{p}_{R}\boldsymbol{p}_{i}}\|\|\overrightarrow{\boldsymbol{p}_{R}\boldsymbol{p}_{g}}\|}>0,\\ -1,&\text{otherwise},\end{cases}$ (7)
@@
where $w_{\text{head}}$, $w_{\text{vel}}$, and $w_{\text{pos}}$ are the weights for each score component. If $S_{i}$ exceeds a threshold $\tau_{\text{leader}}$, we say that Human $i$ is a leader candidate who can lead the robot towards its goal at a desired speed. We add an adjustment term to the previous leader $S_{H_{L}^{t-1}}$ to avoid fluctuation between candidates with similar scores. The candidate with the highest score is selected as $H_{L}^{t}$.
@@ III-B Subgoal Selection
We select the position that is furthest from neighbouring humans:
@@
$v_{\text{max}}=\begin{cases}\|\boldsymbol{v}_{H_{L}^{t}}\|,&\text{if }\|\overrightarrow{\boldsymbol{p}_{R}\boldsymbol{p}_{H_{L}^{t}}}\|\leq\tau_{\text{catchup}},\\ v_{\text{catchup}},&\text{otherwise},\end{cases}$ (12)
where $v_{\text{catchup}}$ is a slightly faster speed when the robot needs to catch up with the leader. For other planners with different parameter settings, this adjustment can be adapted accordingly.
@@ IV Experiments
Figure 4: The three test scenes from the SDD dataset: Promenade, Crossing, and Roundabout. The initial and goal positions are shown in red stars. The direction of travel is interchangeable.
@@ IV-A Simulation Settings
Scenarios. We first perform experiments in simulation environments with real-world trajectory data from the Stanford Drone Dataset (SDD) [39]. SDD was collected mainly from three scenes and we name them Promenade, Crossing, and Roundabout. As shown in Fig. 4, the three scenes exhibit increasing levels of difficulty: 1) Promenade has one main road with side trails, 2) Crossing has a junction with complex traffic, 3) Roundabout involves challenging street crossing. For each scene, we select 10 crowded segments from the longest video. The scene densities are listed in Table II. We evaluate the performance of our method on all 30 segments and use the densest segment from each scene for comparison study. As we directly playback the dataset, the humans are non-reactive to the robot. Reactive human behaviours are studied in the real-world experiments in Sec. IV-D.
[…]
@@
Method Current human states Future human states Environment information Training
@@
Baselines. We compare our method with five crowd navigation methods under different frameworks:
@@
    Social Force (SF) [9] calculates attraction force from the goal and repulsion force from obstacles and other agents.
@@
    HEIGHT [42] is an end-to-end method that is based on deep reinforcement learning and considers crowd interactions and environmental constraints. We follow the training procedures from the provided repository to train the model for holonomic robots.
Evaluation Metrics. We perform 100 repeated runs for each segment and evaluate the average performance to count for the uncertainty of physics engine approximation in the simulation. We mark it as a collision if the robot enters within the agent dimensions. The following three metrics are used for evaluation:
@@
    Total Collision Count (TCC) is the number of frames where collision occurs. We allow the experiment to continue after a collision so there could be multiple counts for one run.

### v2

@@ Abstract
Index Terms:
Human-Aware Motion Planning, Safety in HRI, Social HRI
@@ I Introduction
Crowd navigation is a challenging problem in robotics research as it involves not only static environmental obstacles, but also dynamic agents such as humans. To handle such complex scenarios, recent research focuses on improving the performance of advanced planners with techniques such as predictive planning and reinforcement learning [1, 2, 3, 4].These works aim to develop intelligent and sophisticated robotic systems that can make correct planning decisions, especially in dense crowds and safety-critical scenarios. On the other hand, humans, perhaps the most intelligent agents in the scene, are solving similar navigation problems as the robot. If we think outside the box, can the robot take advantage of human intelligence to simplify its task?
@@
Inspired by how children follow their parents when walking in crowded streets, we find that following trustworthy leaders is an alternative to planning independently, as illustrated in Fig. 1. We can decompose the crowd navigation problem into high-level decisions and low-level short-horizon movements. The former can be solved using the aforementioned intelligent systems or humans, while the latter only requires simple local planners. Therefore, we propose to rethink human-robot relationships in crowd navigation by handing over high-level decisions to humans. We use people as intelligent planners that can help the robot solve the more difficult part of the crowd navigation problem.
@@
Our proposed method consists of two steps, leader selection and subgoal selection. We start by evaluating nearby humans based on their observed trajectories, assuming humans are rational and move consistently within a small time window. Through a set of rule-based criteria, we aim to identify suitable human leaders who are 1) ahead of and reachable from the robot’s current position, 2) heading towards a similar goal at a desired speed. We then find a subgoal such that it allows the robot to follow the selected leader closely with minimal interruption from other humans. The subgoals are designed to be easily achievable using a simple base planner and we choose Social Force (SF) [5] in our experiments. The leader selection process is performed continuously so the robot can switch between different humans for maximised efficiency. We therefore simplify the task from complex crowd navigation to following suitable human leaders using lightweight local planners, which also enables the robot to demonstrate human-like planning behaviours.
Of relevance to our work is the literature on human-following robots. Many of these works study how to maintain a proper distance from humans in different scenarios to ensure human comfort [6, 7].Others study how to identify a specific user in complex environments [8, 9, 10]. In these works, a specific user is predefined for the robot to follow without additional context or objectives. Our work is fundamentally different as 1) the robot actively chooses a suitable leader to follow, 2) by following humans, we effectively address the challenge of crowd navigation.
Our contributions are threefold. First, we propose a novel crowd navigation framework that utilises human intelligence to decompose the original problem into straightforward subgoal planning. Second, we design a set of evaluation processes to find subgoals near suitable human leaders, which can be effectively reached using a simple base planner. Third, we demonstrate through experiments that the proposed People-as-Planner scheme contributes to safe and efficient robot behaviours in crowded and safety-critical scenarios.
@@ II-A Social Robot Navigation
Robot navigation in crowds, or social navigation, has been studied for decades [1, 11]. Traditional model-based methods [12, 5, 13] have been widely adopted in simulations. However, they are less suitable for real-world complex scenarios as they are short-sighted and tend to over-simplify human behaviours [14]. Predictive planning enables the robot to look into the future and avoid potential collisions. They rely heavily on the performance of prediction modules, as inaccurate predictions may lead to accidents and uncertain predictions result in over-conservative behaviours [15, 16]. End-to-end learning-based methods train robot policies using large-scale demonstrations [17, 18, 19].The performance of the trained networks relies on the accuracy of the simulation platform, or the quality of the collected data [20]. Velocity-field-based methods use crowd motion patterns to inform implicit social norms. Their experiment scenarios are usually simple crowd flows in simulations or indoor settings [21, 22, 23, 24], as complex human movements in the real world are difficult to describe using such homogenised fields. Existing methods require the robot to plan its path independently. We propose a different approach where the more complex navigation decisions are solved by humans, and the robot simply needs to reach nearby subgoals.
@@ II-B Human-Following Robot
Examples of social navigation scenarios include nursing and warehouse assistance, where robots are tasked to follow a target user. Such systems generally consist of modules including human detection, human tracking, and path planning [25], which have been extensively studied. Recent works focus on more specific tasks such as frontal following [6], sensor fusion [8, 9, 26], and new application scenarios [27]. In terms of following strategy, some works replicate human trajectories exactly [8, 9], while others define subgoals near the leader [26, 25]. We follow the latter as it reduces the perceived intentionality of tracking and results in more natural behaviours. We direct the reader to [25] for a comprehensive review. Existing human-following methods assign a hypothetical task to the robot to follow a predefined human target. In our work, the robot’s objective is instead to navigate to its destination by actively selecting and following human leaders.
@@ II-C Human-Robot Relationships
As the robot navigates in crowds, there exist complex relationships between humans and the robot. Apart from considering humans as dynamic obstacles, [28] first proposes the concept of using people as sensors to benefit from human intelligence. This line of work infers the positions of occluded obstacles from observed human behaviours [29, 30, 31, 32].Inspired by these works, we propose to use people as planners in crowd navigation. There are a few works that also follow human behaviours in robot planning but with different focuses. [33] and [34] consider human-following as one of the policies when generating trajectory samples in a traditional sampling-based planner. [8] and [35] utilise human guidance for pathfinding in mapless environments with low human density. [36] imitates pedestrian behaviours on a narrow sidewalk with limited forms of interactions between humans and the robot. In comparison, our proposed People-as-Planner framework can be applied to various crowded navigation scenarios and is more suitable for real-world robot deployment.
@@ III Methodology
Problem Definition: The state of an agent consists of its position and velocity $\boldsymbol{x}=[\boldsymbol{p}^{\top},\boldsymbol{v}^{\top}]^{\top}\in\mathbb{R}^{4}$. We use subscripts $H,R$ to represent human and robot, and superscript $t$ to represent timesteps. We assign a unique ID $H_{i}$ where $i\in\{1,\dots,n\}$ to $n$ humans that are within the robot’s observable range. The robot’s goal position is defined as $\boldsymbol{p}_{g}\in\mathds{R}^{2}$. $\boldsymbol{X}_{H}^{0:t}$ represents the historical trajectories of all humans. $\mathcal{O}$ represents the environmental obstacles within the robot’s observable range. In a crowd navigation setting, the robot plan is generated by a planner $\mathcal{P}$, denoted as
@@
The planner $\mathcal{P}$ replans at every step until goal is reached.
Overview: In our proposed scheme, we first identify a human leader $H_{L}^{t}$ and subsequently define a subgoal $\boldsymbol{p}_{g}^{t}$:
@@
where $f_{\text{leader}}(\cdot)$ and $f_{\text{subgoal}}(\cdot)$ are the leader and subgoal selection processes that will be detailed in Sec. III-A and Sec. III-B. Robot plan is then generated by base planner $\mathcal{P}$:
@@
and the process repeats until goal is reached. We assume that the robot dynamics is incorporated into the planner and our method is planner-agnostic. Since each subgoal $\boldsymbol{p}_{g}^{t}$ is much easier to reach than the original goal $\boldsymbol{p}_{g}$, we can use a simple local planner as the base planner. The overall framework is illustrated in Fig. 2.
@@ III-A Leader Selection
Figure 3: An example of reachability check from robot (red cross) to three candidate leaders (green dot). The purple curves enclose the visible region. Human 4 has a larger score than Human 0 for being closer to the robot. Human 3 has a negative score as it is blocked by Human 2.
Next, we evaluate the reachability of each human. As we aim to use simple base planners, the selected leader needs to be directly accessible without the need for complex obstacle avoidance. Spatio-temporal planning [37, 38] is a feasible method, but at a high computational cost with prediction uncertainties. Instead, we adopt Line-of-Sight (LoS) distance metric proposed by [39] to define the reachability. As illustrated in Fig. 3, LoS measures the distance from each human to the boundary of the robot’s visible region constructed from LiDAR scans. We represent the reachability score as $f_{\text{reach}}$ and only consider Human $i$ as a potential leader if
@@
[…]
where $\tilde{\boldsymbol{v}}_{i}=\frac{1}{T}\sum_{k=t-T+1}^{t}\boldsymbol{v}_{i}^{k}$. We choose not to predict future headings because existing prediction methods may overlook corner cases [40] and increase computational costs.
@@
[…]
@@
Third, we compare relative positions to identify humans between robot and its goal, favouring the closer ones:
[…]
@@
where $w_{\text{head}}$, $w_{\text{vel}}$, and $w_{\text{pos}}$ are the weights for each score component. If $S_{i}$ exceeds a threshold $\tau_{\text{leader}}$, Human $i$ becomes a leader candidate who can lead the robot towards its goal at a desired speed. We add an adjustment term to the previous leader $S_{H_{L}^{t-1}}$ to avoid fluctuation between candidates with similar scores. The candidate with the highest score is selected as $H_{L}^{t}$.
@@ III-B Subgoal Selection
We select the position furthest from neighbouring humans:
@@
[…]
where $v_{\text{catchup}}$ is a slightly faster speed when the robot needs to catch up with the leader. For other planners with different settings, this adjustment can be adapted accordingly.
@@ IV Experiments
Figure 4: Three test scenes from the SDD dataset: Promenade, Crossing, and Roundabout. The initial and goal positions are shown in red stars. The direction of travel is interchangeable.
@@ IV-A Simulation Settings
Scenarios: We first perform simulation experiments with real-world trajectory data from the Stanford Drone Dataset (SDD) [41]. SDD was collected mainly from three scenes and we name them Promenade, Crossing, and Roundabout. As shown in Fig. 4, the three scenes exhibit increasing levels of difficulty: 1) Promenade has one main road, 2) Crossing has a junction with complex traffic, 3) Roundabout involves challenging street crossing. For each scene, we select 10 crowded segments from the longest video. The scene densities are listed in Table II. We evaluate our method on all 30 segments and use the densest segment from each scene for comparison study. As we directly playback the dataset, the humans are non-reactive. Reactive human behaviours are studied in the real-world experiments in Sec. IV-D.
Implementation Details: We use a holonomic robot model and the control dynamics is incorporated into the base SF planner. We set $v_{\text{pref}}$, $r$, $d$ to $1.4\text{m/s}$ [42], $10\text{m}$, and $0.8\text{m}$ respectively. We set the robot speed limit to $2.0\text{m/s}$. The rest of the tunable parameters can be found in our code repository. We find the set of parameters through manual tuning and we use the same values for all experiments. Since not all baselines distinguish heterogeneous road users, we set the radius of all agents, including robot, humans, bicycles, and cars to 0.5m for fair comparisons. For collision evaluation, we provide two sets of agent dimensions: 1) all agents have the same radius of 0.5m, 2) bicycles and vehicles are given realistic dimensions of $1.9\text{m}\times 1\text{m}$ and $4.5\text{m}\times 1.9\text{m}$. We report results for both settings in Sec. IV-B. Simulations are conducted at 30Hz on Gazebo platform [43] with AMD 5950XT CPU and RTX3090 graphic card.
@@
Method
Current
human states
Future
human states
Environment
information
Training
@@
Baselines: We compare our method with five crowd navigation methods under different frameworks:
@@
    Social Force (SF) [5] calculates attraction from goal and repulsion from obstacles and neighbours.
@@
    HEIGHT [44] is an end-to-end reinforcement learning method that considers crowd interactions. We follow the training procedures from the provided repository to train the model for holonomic robots.
Evaluation Metrics: We perform 100 repeated runs for each segment and evaluate the average performance to count for the uncertainty of physics engine approximation in simulation. We mark it as a collision if the robot enters within the agent dimensions. The following three metrics are used:
@@
    Total Collision Count (TCC) is the number of frames where collision occurs. Experiment continues after a collision so there could be multiple counts in one run.